OpenStudy (anonymous):

Evaluate f(x) = 5x + 3, if x = -4. (Points : 2) 23 -17 17 -23

OpenStudy (anonymous):

-17

OpenStudy (adi3):

do u know whts f(x) is

OpenStudy (anonymous):

Thats what I thought. Just wanted to make sure. To prove it. Ill show how

OpenStudy (adi3):

5(-4) + 3 =

OpenStudy (adi3):

@DelTa-_-LeGEnD

OpenStudy (anonymous):

5*-4= -20, -20+3= -17

OpenStudy (anonymous):

F(x) Plug in -4 for X So then you have 5(-4) +3= -17

OpenStudy (anonymous):

A line with undefined slope is a function. (Points : 1) True False

OpenStudy (anonymous):

My guess is false

OpenStudy (adi3):

true

OpenStudy (anonymous):

The answer is true? How?

OpenStudy (adi3):

u can solve it by looking at where is the line

OpenStudy (anonymous):

?

OpenStudy (adi3):

trust me check ur answer its true

OpenStudy (anonymous):

What is the direct variation equation if f(x) varies directly with x and f(x) = –9 when x = 3? A. f(x) = -1/3x B. f(x)= -3x C. f(x)= -6x D.f(x)= -12x

OpenStudy (anonymous):

@Adi3

OpenStudy (adi3):

wht

OpenStudy (adi3):

quick i hve to sleep

OpenStudy (adi3):

its 11:00 P.M

OpenStudy (texaschic101):

a line with an undefined slope is not a function....an undefined slope represents a vertical line where x is always the same number....and a function can not have 2 or more repeating x values

OpenStudy (anonymous):

I posted an new one above

OpenStudy (adi3):

-6

OpenStudy (adi3):

u gonna medal me

OpenStudy (anonymous):

There :)

OpenStudy (anonymous):

What is the vertex of f(x) = |x + 3| - 5 ? Your answer should be an ordered pair.